# Understanding Ethical Practices in AI: Insights from a Cross-Role, Cross-Region Survey of AI Development Teams

### v1

@@ 5.4.2. RA Role - Mitigation Strategies
We examined which AI ethics principles RAs prioritize when defining AI requirements and found notable variation by geographic region. In North America, 66.67% of participants, and in the EU+EEA+UK, 88.89% of participants, identified “data protection and the right to privacy” as their top priority. In contrast, participants from other regions ranked this third, behind “democracy and rule of law” (62.50%) and “harm prevention and beneficence” (50%). “Transparency and explainability of AI systems” ranked second among RAs in both North America (53.33%) and the EU+EEA+UK (55.56%), but dropped to fifth place among RAs from other regions (25%). These differences in prioritization may reflect the influence of existing regulatory frameworks on developers’ ethical considerations. RAs in North America and the EU+EEA+UK tended to emphasize principles, such as “data protection and the right to privacy” and “transparency and explainability of AI systems,” that align closely with established governance frameworks, including the EU’s General Data Protection Regulation (GDPR) (Commission, 2018), the California Consumer Privacy Act (CCPA) (Government of California, 2022), the EU AI Act (Commission, 2024). By contrast, RAs from other regions emphasized principles, such as “democracy and rule of law” and “harm prevention and beneficence,” which align more with universal ethical standards than specific regional regulations.
Table 13. Frequency of Ethical Consideration in AI Requirements Documentation
Always Often Sometimes Rarely Never
Count 7 12 7 6 2
Percentage 20.6% 35.3% 20.6% 17.6% 5.9%
Table 14. AR Roles: Impacts of Ethics Requirements on AI Projects
@@
Our results indicate that teams with AM and AR role members who are more familiar with AI ethics principles are more likely to integrate these principles during both documentation and design. Among the most common strategies AM roles employ are the establishment of ethical guidelines and effective communication with team members. Overall, AM roles could increase their impact by delegating greater ethical decision-making authority to compliance teams and establishing dedicated ethics councils.
5.4.3. AD Role - Mitigation Strategies
@@
Figure 11. AD Roles: Methods for Bias Mitigation
@@
The majority of participants in AD roles (76.2%) consider transparency and explainability of AI systems at least “moderately important,” while only about 8.7% consider it either “not at all important” or “not very important” (see Table 15). This aligns with our findings on familiarity with AI ethics principles, where AD roles reported being, on average, very familiar with this principle (see Figure 5 for more details). Similarly, a large portion of participants (64.1%) are at least “moderately confident” in their ethical decisions during AI development. A small number (12.2%) appears to be not very confident.
Table 15. AD Roles: Importance of Transparency and Explainability of AI Systems
Response Option Count Percentage
Extremely Important 73 52.5%
Moderately Important 33 23.7%
Somewhat Important 21 15.1%
Not Very Important 7 5.1%
Not At All Important 5 3.6%
Table 16. AD Roles: Confidence in Making Ethical Decisions During AI Development.
Response Option Count Percentage
Extremely Confident 35 25.2%
Moderately Confident 54 38.9%
Somewhat Confident 32 23.0%
Not Very Confident 11 7.9%
Not At All Confident 6 4.3%
Prefer not to say 1 0.7%
@@
We then evaluated whether there is a correlation between participants in AD roles’ average familiarity with AI ethics principles and two factors: the importance they assign to AI systems being transparent and explainable (H4c), and their confidence in making ethical decisions during AI development (H4d) (see Table 10). With a pvalue<0.0001p-value<0.0001 and $\chi^{2}=45.8273$ for H4c, we can reject the null hypothesis and conclude that there is a correlation between the average familiarity with our AI ethics principles and the importance of transparent and explainable AI systems. With a pvalue=0.0050p-value=0.0050 for H4d, we cannot reject the null hypothesis after applying the Bonferroni correction. However, this value is below the conventional alpha threshold of 0.05, which could suggest the presence of a correlation, though not a statistically significant one.
@@
We also aimed to identify what types of training or resources AD roles can use to help them better integrate AI ethics. Figure 12 shows the result. “Access to ethical guidelines and best practices” and “collaboration with AI ethicists and legal experts” ranked the highest (59.85% and 55.47%, respectively). Interestingly, “support for continuous education related to AI ethics” appears to be the least popular resource (39.42%).
@@
To gain insights into the individual practices of AD roles, we asked this group of participants to describe how they integrate AI ethics principles into their coding practices and algorithm designs. Among the diverse set of 115 answers, we found several themes. 22 ADs described methods completed during the data processing phases (e.g.,“I ensure that the data used to train AI models is diverse and representative to prevent bias. I also use techniques like re-sampling, data augmentation, or bias-correction algorithms to minimize any inherent biases in the data”). We also observed that data processing is essential for mitigating biases in AI systems today, as one AD mentioned: “I ensure that the data used to train AI models is diverse and representative to prevent bias. I also use techniques like re-sampling, data augmentation, or bias-correction algorithms to minimize any inherent biases in the data.” Another participant replied: “When collecting data, ensure the legitimacy and compliance of the data source and avoid violating user privacy.” Yet another common approach for ADs is establishing and adhering to AI guidelines. These may coincide with the education and training for developers, as one participant noted: “We have take mandatory courses on AI ethics. There are also somewhat frequent workshops on AI ethics. Also try to align the system along existing frameworks on AI ethics.” These guidelines may take inspiration from the regulatory framework as one AD responded: “maintaining defined protocols and regulations.”
@@
Similar to RA roles, we asked participants in AD roles if their company conducts (ethics-related) risk and impact assessments. Of the 107 responses from ADs, 64.5% indicated “yes,” while 24.3% and 6.5% indicated “no” or were “unsure,” respectively. A small number preferred not to respond. The positive responses are slightly higher than those from RA roles, suggesting that both roles are actively involved in risk assessments and ensuring ethical guidelines are considered during the requirements and development phases. There is a variety in how participants understood and described these activities, with one saying: “The security team in my company conduct regular security and compliance assessments of all of our software that are shipped to production.” One participant noted, “Peer reviews,” while another added: “Yes, we do, by looking at previous case studies and possible pitfalls to make sure we stay within parameters.”
Figure 12. AD Roles: Training Resources to Help with AI Ethics
5.4.4. QA Role - Mitigation Strategies
@@
For participants in QA roles, we aimed to understand which techniques they use to identify algorithmic biases. As shown in Figure 13, the top strategy is “identifying and testing results for vulnerable groups in AI systems” (66.67%). Interestingly, this approach is among the least commonly selected by participants in AD roles (see Figure 11). “User testing with diverse groups” is the second most common strategy (60%). The rest of the strategies are loosely aligned with those chosen by AD roles, focusing on checking the training data and evaluating the model output. “Evaluating correct implementation of fairness constraints in the model” (20%) is given similarly low priority by QA roles as it is by AD roles (i.e., “implementing fairness constraints in the model”) (33.33%).
@@
We hypothesized that a positive relationship exists between QA participants’ average familiarity with AI ethics principles and the importance they attribute to AI ethics in their role. (H4e). With pvalues=0.8270p-values=0.8270, we cannot reject the null hypothesis and see no statistical relationship between QA’s familiarity with principles and their perceived importance of AI ethics training.
Figure 13. QA Roles: Techniques to Identify Biases in AI Systems
Table 17. QA Roles: Importance of Training on AI Ethics for QA or Tester Role
Response Option Count Percentage
Extremely Important 6 40.00%
Somewhat Important 5 33.33%
Moderately Important 2 13.33%
Not Very Important 2 13.33%
Figure 14. QA Role: How Data Privacy is Maintained in AI Testing
@@
Thirteen participants in QA roles described various steps they take to ensure that AI systems comply with ethical guidelines and prevent biased outputs. One participant responded: “We user test with a large, diverse sample. Also, we are constantly updating and enhancing diversity training.” This approach aligns with the importance of diverse voices in development teams for minimizing biases, emphasized in previous research, such as (Kawakami et al., 2024; Olson et al., 2025). Other participants provided insights into their team’s internal processes, with one responding: “We have already got data which can be used internally and there is a team which regulates this part of testing. hence, we take care of this.” Another QA responded “I always follow data privacy policies at regular interval. My first duty is to follow and use data security policies.”
@@
Furthermore, we asked participants about their approaches to ensure data privacy is maintained during AI testing. Most of them (80%) mentioned “using secure testing environment,” while other privacy-related techniques, such as “test data anonymization” and “conducting privacy impact assessment (PIA),” were also considered. PIA is generally required to be conducted prior to planning data collection and not during the QA and testing phases. However, consistent with (Prybylo et al., 2024), we observed that PIA is often performed at other stages of development. Lastly, “using synthetic data” was not a commonly adopted method (see Figure 14).
@@
Similar to RA and AD roles, we asked QAs if their company conducts (ethics-related) risk and impact assessments. Of the thirteen responses, eight said “yes.” However, their perspectives varied, with one saying, “We need to comply with EU laws. It is needed to be checked if our app doesn’t suggest wrong products or suggests something wrong to the user.” Another participant added: “This was done as part of the initial part, now things are streamlined,” while another said: “No, my company does not conduct risk and impact assessments or human rights due diligence. This is because we currently lack the resources and processes needed to carry out these evaluations effectively.”
5.4.5. ISec Role - Mitigation Strategies
@@
ISec members from Group B had varied perspectives on the impact of AI technologies on data privacy within their company. When asked about their perspectives on this impact, 46.14% chose “somewhat” or “significantly compromise data privacy,” while 38.26% noted they “somewhat” or “significantly improve data privacy.” The remaining 15.38% said there is “no impact on data privacy at all.” Table 18 presents these results.
@@
Table 18. ISec Roles: Impact of AI Technologies on Data Privacy within the Company
@@
When asked what measures their company implements to mitigate AI-related security or privacy risks (see Figure 15), traditional security measures, such as “access controls and user authentication” and “encryption of data used in AI models,” are the top two approaches (91.67% and 58.33%, respectively). This result aligns with prior research suggesting that security concepts are generally more tangible than privacy concepts (Hadar et al., 2018; Prybylo et al., 2024). Half of the participants consider implementing “incident response plans specific to AI security/privacy breaches,” or conducting “regular security and/or privacy audits of AI systems.” Only 33.33% of the participants said they “integrate privacy-enhanced technologies” or “collaborate with AI developers to identify security/privacy risks.” This suggests a potential need for ISec members to collaborate more closely with developers while implementing incident response plans and conducting AI audits.
Figure 15. ISec Roles: Measures to Mitigate AI-Related Security and Privacy Concerns
@@
We asked ISec members how they address potential ethical concerns about data privacy and security in AI systems. From twelve responses, we observed a variety of approaches, with five participants mentioning the use of various security protocols. For instance, one participant responded: “To address ethical concerns related to data privacy and security in AI systems, I implement strong data protection measures, ensure compliance with privacy regulations, use encryption and access controls, and regularly audit and update security practices. Additionally, I advocate for transparency and user consent regarding data usage.” However, three of these five responses limited their scope to only traditional security methods (e.g., “Use encryption, access controls, and other security measures to protect data from unauthorized access”). Another group member noted: “I try to use common sense, but It is handled by other areas.”
@@
ISec members were then asked how they or their organizations ensure the protection of personal data and uphold the privacy rights of affected communities. Six out of the eleven responses mentioned traditional security or privacy techniques, e.g., “By introducing encryption to every system” or “Collect only the necessary data for the intended purpose and avoid excessive data collection.” Three of the eleven responses also mention following policies or guidelines, with one ISec member responding: “Yes, we have AI policy in place that is strictly adhere by all employees. We also have configured DLP tools that help detect misuse of PII data.” Another participant simply said: “we follow main GDPR recommandations.” ISec members’ responses highlight the importance of following standard data security and privacy practices in AI development and the need for improvement due to a lack of awareness around emerging ethical concerns associated with AI systems.
@@
Table 19. ISec Roles: Involvement in Integrating Privacy and Security Principles into AI Development
Response Option Count Percentage
Somewhat Involved 5 38.46%
Very Involved 4 30.77%
Rarely Involved 2 15.38%
Neutral 1 7.69%
Prefer not to say 1 7.69%
@@
AD and QA roles employ complementary methods to mitigate biases in models, results, and training data. For ADs, access to ethical guidelines and best practices, along with collaboration with AI ethicists and legal experts, is highly valued. QA team members, meanwhile, consider AI ethics training to be at least somewhat important and focus primarily on maintaining secure testing environments, anonymizing test data, and conducting privacy impact assessments (PIAs) to protect data privacy during testing. In contrast, those in ISec roles tend to rely more heavily on traditional security measures such as access control, authentication, and encryption. QA and AD professionals are generally less involved in ethics-related impact assessments than those in AM, RA, and ISec roles. Improving communication and collaboration across development teams could support a more holistic and integrated approach to AI ethics.
5.4.6. AR + AE Role - Mitigation Strategies

### v2

@@ 4.4.2 Requirements Analyst (RA) Roles
We examined which AI ethics principles RAs prioritize when defining AI requirements and found notable variation by geographic region. In North America and the EU+EEA+UK, 66.67% and 88.89% of participants identified “data protection and the right to privacy” as their top priority. In contrast, participants from other regions ranked this principle third, behind “democracy and rule of law” (62.50%) and “harm prevention and beneficence” (50%). “Transparency and explainability of AI systems” ranked second among RAs in both North America (53.33%) and the EU+EEA+UK (55.56%), but dropped to fifth place among RAs from other regions (25%). These differences in prioritization may reflect the influence of existing regulatory frameworks on developers’ ethical considerations. RAs in North America and the EU+EEA+UK tended to emphasize principles, such as “data protection and the right to privacy” and “transparency and explainability of AI systems,” that align closely with established governance frameworks, including the EU’s General Data Protection Regulation (GDPR) [GDPR2018], the California Consumer Privacy Act (CCPA) [CCPA], the EU AI Act [AIAct2024]. By contrast, RAs from other regions emphasized principles, such as “democracy and rule of law” and “harm prevention and beneficence,” which align more with universal ethical standards than specific regional regulations.
Table 20: RA Roles: Impacts of Ethics Requirements on AI Projects
@@
RQ4 - AM & RA Roles: Our results reveal that those in AM and RA roles who are more familiar with AI ethics principles are more likely to integrate them during both documentation and design. Among the most common strategies they employ are establishing ethical guidelines and maintaining effective communication with team members. Overall, AM roles could increase their impact by delegating greater ethical decision-making authority to compliance teams and creating dedicated ethics councils.
4.4.3 AI Developer, Architect, and Data Scientist (AD) Roles
@@
Figure 10: AD Roles: Methods for Bias Mitigation
@@
The majority of those in AD roles (76.2%) consider transparency and explainability of AI systems at least “moderately important,” while only about 8.7% consider it either “not at all important” or “not very important” (see Table 21). This aligns with our findings on familiarity with AI ethics principles, where AD roles, on average, are very familiar with this principle (see Figure 5 for more details). Similarly, a large portion of participants (64.1%) are at least “moderately confident” in their ethical decisions during AI development. A small number (12.2%) appears to be not very confident.
Table 21: AD Roles: Importance of Transparency and Explainability of AI Systems
Extremely Important Moderately Important Somewhat Important Not Very Important Not At All Important
52.5% 23.7% 15.1% 5.1% 3.6%
@@
We then evaluated whether there is a correlation between AD roles’ average familiarity with AI ethics principles and two factors: the importance they assign to AI systems being transparent and explainable (H5c), and their confidence in making ethical decisions during AI development (H5d) (see Table 16). With a pvalue<0.0001p-value<0.0001 and $\chi^{2}=45.8273$ for H5c, we can reject the null hypothesis and conclude that there is a correlation between the average familiarity with AI ethics principles and the importance of transparent and explainable AI systems. With a pvalue=0.0050p-value=0.0050 for H5d, we cannot reject the null hypothesis after applying the Bonferroni correction. However, this value is below the conventional alpha threshold of 0.05, which could suggest the presence of a correlation, though not a statistically significant one.
@@
We analyzed what types of training or resources AD roles may use to help them better integrate AI ethics. Figure 11 shows the result. “Access to ethical guidelines and best practices” and “collaboration with AI ethicists and legal experts” ranked the highest (59.85% and 55.47%, respectively). Interestingly, “support for continuous education related to AI ethics” appears to be the least popular resource (39.42%).
@@
To gain insights into the individual practices of AD roles, we asked them to describe how they integrate AI ethics principles into their coding practices and algorithm designs. Among the diverse set of 115 answers, we found several themes. 22 ADs described methods completed during the data processing phases (e.g.,“I ensure that the data used to train AI models is diverse and representative to prevent bias. I also use techniques like re-sampling, data augmentation, or bias-correction algorithms to minimize any inherent biases in the data”). We also observed that data processing is essential for mitigating biases in AI systems today, as one AD mentioned: “I ensure that the data used to train AI models is diverse and representative to prevent bias. I also use techniques like re-sampling, data augmentation, or bias-correction algorithms to minimize any inherent biases in the data.” Another participant replied: “When collecting data, ensure the legitimacy and compliance of the data source and avoid violating user privacy.” Yet another common approach is establishing and adhering to AI guidelines. These may coincide with the education and training for developers, as one noted: “We have take mandatory courses on AI ethics. There are also somewhat frequent workshops on AI ethics. Also try to align the system along existing frameworks on AI ethics.” These guidelines may take inspiration from the regulatory framework as one AD responded: “maintaining defined protocols and regulations.”
@@
Similar to RA roles, we asked those in AD roles if their company conducts (ethics-related) risk and impact assessments. Of the 107 responses, 64.5% indicated “yes,” while 24.3% and 6.5% indicated “no” or were “unsure,” respectively. A small number preferred not to respond. The positive responses are slightly higher than those from RA roles, suggesting that both roles are actively involved in risk assessments and ensuring ethical guidelines are considered during the requirements and development phases. There is a variety in how they understood and described these activities, with one saying: “The security team in my company conduct regular security and compliance assessments of all of our software that are shipped to production.” One participant noted, “Peer reviews,” while another added: “Yes, we do, by looking at previous case studies and possible pitfalls to make sure we stay within parameters.”
Table 22: AD Roles: Confidence in Making Ethical Decisions During AI Development
Extremely Confident Moderately Confident Somewhat Confident Not Very Confident Not Confident At All Prefer not to say
25.2% 38.9% 23.0% 7.9% 4.3% 0.7%
Figure 11: AD Roles: Training Resources to Help with AI Ethics
4.4.4 Quality Assurance Engineer and IT Maintenance (QA) Roles
@@
We aimed to understand which techniques those in QA roles use to identify algorithmic biases. As shown in Figure 12, the top strategy is “identifying and testing results for vulnerable groups in AI systems” (66.67%). Interestingly, this approach is among the least commonly selected by participants in AD roles (see Figure 10). “User testing with diverse groups” is the second most common strategy (60%). The remaining strategies are loosely aligned with those chosen by AD roles, focusing on checking the training data and evaluating the model output. “Evaluating correct implementation of fairness constraints in the model” (20%) is given similarly low priority by QA roles as by AD roles (i.e., “implementing fairness constraints in the model”) (33.33%).
@@
We hypothesized that a positive relationship exists between the average familiarity with AI ethics principles of those in QA roles and the importance they attribute to AI ethics in their role (H5e). With pvalues=0.8270p-values=0.8270, we cannot reject the null hypothesis and see no statistical relationship between QA’s familiarity with principles and their perceived importance of AI ethics training.
Figure 12: QA Roles: Techniques to Identify Biases in AI Systems
Table 23: QA Roles: Importance of Training on AI Ethics for QA or Tester Role
Extremely Important Moderately Important Somewhat Important Not Very Important
40.00% 33.33% 13.33% 13.33%
Figure 13: QA Roles: How Data Privacy is Maintained in AI Testing
@@
Thirteen participants in QA roles described various steps they take to ensure that AI systems comply with ethical guidelines and prevent biased outputs. One participant responded: “We user test with a large, diverse sample. Also, we are constantly updating and enhancing diversity training.” This approach aligns with the importance of diverse voices in development teams for minimizing biases, emphasized in previous research, such as [kawakami2024studyingup, olson2025speaks]. Other participants provided insights into their team’s internal processes, with one responding: “We have already got data which can be used internally and there is a team which regulates this part of testing. hence, we take care of this.” Another QA responded “I always follow data privacy policies at regular interval. My first duty is to follow and use data security policies.”
@@
Furthermore, we asked them about their approaches to ensure data privacy is maintained during AI testing. Most of them (80%) mentioned “using secure testing environment,” while other privacy-related techniques, such as “test data anonymization” and “conducting privacy impact assessment (PIA),” were also considered. PIA is generally required to be conducted prior to planning data collection, rather than during the QA and testing phases. However, consistent with findings in [prybylo2024evaluating], we observed that PIA is often performed at other stages of development. Lastly, “using synthetic data” was not a commonly adopted method (see Figure 13).
@@
Similar to RA and AD roles, we asked QAs if their company conducts (ethics-related) risk and impact assessments. Of the thirteen, eight said “yes.” However, their perspectives varied, with one saying, “We need to comply with EU laws. It is needed to be checked if our app doesn’t suggest wrong products or suggests something wrong to the user.” Another participant added: “This was done as part of the initial part, now things are streamlined,” while another said: “No, my company does not conduct risk and impact assessments or human rights due diligence. This is because we currently lack the resources and processes needed to carry out these evaluations effectively.”
4.4.5 Information Security and Privacy Expert (ISec) Roles
@@
ISec members had varied perspectives on the impact of AI technologies on data privacy within their company. 46.14% mentioned that AI “somewhat” or “significantly compromise data privacy,” while 38.26% noted they “somewhat” or “significantly improve data privacy.” The remaining 15.38% said there is “no impact on data privacy at all.” Table 24 presents these results.
@@
Table 24: ISec Roles: Impact of AI on Data Privacy within the Company
@@
When asked what measures their company implements to mitigate AI-related security or privacy risks (see Figure 14), traditional security measures, such as “access controls and user authentication” and “encryption of data used in AI models,” are the top two approaches (91.67% and 58.33%, respectively). This result aligns with prior research suggesting that security concepts are generally more tangible than privacy concepts [hadar2018privacy, prybylo2024evaluating]. Half of them consider implementing “incident response plans specific to AI security/privacy breaches,” or conducting “regular security and/or privacy audits of AI systems.” Only 33.33% said they “integrate privacy-enhanced technologies” or “collaborate with AI developers to identify security/privacy risks.” This suggests a potential need for ISec members to collaborate more closely with developers while implementing incident response plans and conducting AI audits.
Figure 14: ISec Roles: Measures to Mitigate AI-Related Security and Privacy Concerns
@@
ISec members described how they address potential ethical concerns about data privacy and security in AI systems. From twelve responses, we observed a variety of approaches, with five participants mentioning the use of various security protocols. For instance, one participant responded: “To address ethical concerns related to data privacy and security in AI systems, I implement strong data protection measures, ensure compliance with privacy regulations, use encryption and access controls, and regularly audit and update security practices. Additionally, I advocate for transparency and user consent regarding data usage.” However, three of these five responses limited their scope to only traditional security methods (e.g., “Use encryption, access controls, and other security measures to protect data from unauthorized access”). Another group member noted: “I try to use common sense, but It is handled by other areas.”
@@
They also explained how they or their organizations ensure the protection of personal data and uphold the privacy rights of affected communities. Six out of the eleven responses mentioned traditional security or privacy techniques, e.g., “By introducing encryption to every system” or “Collect only the necessary data for the intended purpose and avoid excessive data collection.” Three of them also mention “following policies or guidelines”, with one ISec member responding: “Yes, we have AI policy in place that is strictly adhere by all employees. We also have configured DLP tools that help detect misuse of PII data.” Another participant simply said: “we follow main GDPR recommandations.” These responses highlight the importance of following standard data security and privacy practices in AI development and the need for enhancing awareness around emerging ethical concerns associated with AI systems.
@@
Table 25: ISec Roles: Involvement in Integrating Privacy and Security Principles
Very Involved Somewhat Involved Rarely Involved Neutral Prefer not to say
30.77% 38.46% 15.38% 7.69% 7.69%
@@
RQ4 - AD, QA, & ISec Roles: AD and QA roles employ complementary methods to mitigate biases in models, results, and training data. For ADs, access to ethics guidelines and best practices, along with collaboration with AI ethicists and legal experts, is highly valued. QA members consider AI ethics training to be at least somewhat important and focus primarily on maintaining secure testing environments, anonymizing test data, and conducting PIAs to protect data privacy during testing. In contrast, those in ISec roles tend to rely more heavily on traditional security measures such as access control, authentication, and encryption. QA and AD professionals are generally less involved in ethics-related impact assessments than those in AM, RA, and ISec roles. Enhancing communication and collaboration across development teams may support a more holistic approach to AI ethics.
4.4.6 AI Researcher (AR) and AI Ethicist (AE) Roles